OpenStudy (samigupta8):

Let w be a complex number.Then the set of all complex numbers z satisfying the equation w-w1z =k(1-z) for some real number k Here w1 is conjugate of w...

OpenStudy (samigupta8):

@parthkohli

OpenStudy (samigupta8):

@priyar

OpenStudy (samigupta8):

@faiqraees

OpenStudy (samigupta8):

@phi

OpenStudy (phi):

how far did you get ?

OpenStudy (samigupta8):

I assumed that w is a+ib and z is x+iy... Then i solve the LHS n RHS...as per the equation

OpenStudy (samigupta8):

I got this.. a-ax-by=k(1-x) b+bx-ay=-ky

OpenStudy (phi):

I think you get \[ z = \frac{k-w}{k-w'}\] now I'm trying to figure out what that means (perhaps geometrically?)

OpenStudy (phi):

or z= (w-k)/(w'-k) (a-k) + bi / (a-k) - bi in polar coords, they both have the same magnitude so z will have magnitude 1

OpenStudy (samigupta8):

I m getting it as z=(-1)e^itheta Theta is argument for w

OpenStudy (phi):

I'm would think it's 2 * theta z= exp(i theta) / exp(-i theta) = exp(i 2 theta)

OpenStudy (phi):

wait, my theta is the arg of < (a-k) , b>

OpenStudy (samigupta8):

Yep....U r very correct....@phi ur method of modulus was apt...

OpenStudy (samigupta8):

Same moduli of the numerator n denominator gives us 1

OpenStudy (samigupta8):

Which would serve the modulus of z

OpenStudy (samigupta8):

Thanks...@phi

OpenStudy (phi):

I think as you vary k, we can generate any angle we want, so the solutions will lie on the unit circle. But double check my thinking.